OpenStudy (anonymous):

A rational expression has been simplified below. (x+4)(x-2)/9(x-2) = x+4/9 For what values of x are the two expressions equal?

OpenStudy (danjs):

\[\frac{ (x+4)(x-2) }{ 9*(x-2) } = \frac{ x+4 }{ 9 }\] that look right

OpenStudy (anonymous):

\[all R except x=2\]

OpenStudy (danjs):

yes, th eleft is the same as the right multiplied by 1 or (x-2)/(x-2)

OpenStudy (danjs):

that introduces an extraneous answer, the bottom can not be zero. x can not be 2